OpenStudy (anonymous):

solve using the multiplication principle -1/8x=-9/10

OpenStudy (anonymous):

Ok to solve for x you need to separate the x from the -1/8. you can do that two ways. You can either divide both sides by -1/8, or you can multiply by the inverse fraction (-8/1) to both sides.

OpenStudy (anonymous):

*x, *-9/10 x = 10/72

OpenStudy (anonymous):

*-10/9, I mean.

OpenStudy (anonymous):

That is not what I get, you should be able to multiply both sides by (-8/1) = -8, which gives you x=72/10 = 36/5 = 7 1/5

OpenStudy (anonymous):

so the correct answ is -10/9

OpenStudy (anonymous):

right thts wht i cam up with 10/72

OpenStudy (anonymous):

72/10

OpenStudy (anonymous):

I am not sure what Msc. did

OpenStudy (anonymous):

me either

OpenStudy (anonymous):

I just divied 72/10 which gave me 7.2

OpenStudy (anonymous):

-1/8x = -9/10 => 1/8x=9/10 Multiply by 10: 10/8x=9 Multiply by x: 9x=10/8 Divide by 9: x=10/72

OpenStudy (anonymous):

is tht the answer

OpenStudy (anonymous):

That is the correct answer, yes.

OpenStudy (anonymous):

I just ran it through my calc, it comes up with 7.2 also. I think you made a mistake in the second step, you can't multiple to move the x from one side to the other, you would have had to divide.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

so x=?

OpenStudy (anonymous):

It says 1/8x, not x/8 or (1/8)x. What I wrote is correct.

OpenStudy (anonymous):

10/72 is tht correct

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Technically, you should always provide the simplest solution, so a better answer would be 5/36.

OpenStudy (anonymous):

x = -9/10 x -8 = 72/10 = 7.2

OpenStudy (anonymous):

is this simplified it got to be simplify or fractions

OpenStudy (anonymous):

That is not correct. It clearly says (-1)/8x=(-9)/10. You have to invert the equation, or multiply by x in order to isolate for x. Unless the OP's equation is incorrect. x=5/36 is the correct answer.

OpenStudy (anonymous):

ok thank you so much for being patient with me

OpenStudy (anonymous):

No problem.

OpenStudy (anonymous):

thank all of u

OpenStudy (anonymous):

oh ok - i took the LHS to be (-1/8) * x my mistake

OpenStudy (anonymous):

That's what I figured ;)